y2+yt-12t2 show me how i can work this problem
Answer by Cromlix(4381)

y^2 + yt - 12t^2
(y + 4t)(y - 3t)= 0
y + 4t = 0
y = -4t
y - 3t = 0
y = 3t.
Hope this helps.
:-)